# Cobblestone: A Divide-and-Conquer Approach for Automating Formal Verification

### v3

@@ 2.6. Recursively Invoking Cobblestone
Through this recursion, Cobblestone uses divide-and-conquer, iteratively simplifying the goal and recursing on simpler subgoals, increasing the probability of success. This way, Cobblestone breaks down the problem of proving the original theorem into smaller chunks, and leverage progress made in the previous iterations to assist the subsequent ones. Cobblestone recurses in a greedy depth-first manner; preliminary experiments with backtracking and restarting from scratch do not improve performance.
@@ 3. Evaluation
We evaluate Cobblestone on four datasets of theorems from open-source Coq projects. We compare Cobblestone to five prior tools — CoqHammer, Proverbot9001, Tactician, PALM, and Rango— as well as two LLM baselines of our own creation.
Our evaluation answers five research questions:
@@ 3.1. Experimental Setup
Benchmarks. We construct our evaluation benchmarks from CoqGym’s test set of theorems (Yang19, ), and from three other open-source projects — coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ).
@@
CoqGym is a widely used benchmark for evaluating proof-synthesis tools (Yang19, ; First20oopsla, ; First22icse, ; Sanchez-Stern23toplas, ), comprised of 68,501 theorems and their associated human-written proofs across 124 projects. CoqGym’s test set consists of 26 projects with 12,161 theorems, including verdi-raft (distributed software correctness), goedel (Gödel’s 1st incompleteness theorem), and tree-automata (verified algorithms). Evaluating using this dataset allows for a more fair comparison to prior tools, and assessing tool efficacy across a diverse set of projects. However, CoqGym was released in 2019, and its projects existed on GitHub before then. Because the pre-training cutoff for GPT-4 is September 2021, GPT-4 has potentially seen CoqGym in its pre-training.
@@
To address this concern, we also evaluate using coq-wigderson (10.1145/3618305.3623600, ), coq-bb5 (bb5github, ), and PnVRocqLib (pnvgithub, ). The first commit from each of these projects (March 2022, April 2024, and August 2024, respectively) is after the GPT-4 pre-training cutoff. coq-wigderson is a formal verification of Wigderson’s graph coloring algorithm. It consists of 174 theorems. PnVRocqLib contains 810 theorems about mathematics, including first order logic, Hilbert logic, and boolean algebra. coq-bb5 contains 1446 theorems about busy-beaver values.
@@
Because of the cost of performing inference on LLMs, full-scale evaluations on tens of thousands of theorems are impractical. In fact, the cost of just the LLM usage for our evaluation, including ablation studies, exceeded US\$6K. Instead, we create four 100-proof subsets of the four benchmark projects, sampling theorems at random. We call these subsets CoqGym100, Wigderson100, PnVRocqLib100, and BB5100. To keep the cost of ablations low, we evaluate on PnVRocqLib100 and BB5100 only for RQ1 (Section 3.2). The total size of this dataset (400 theorems), is in line with other prior work which uses LLMs for proof synthesis (wanglego, ; wang2024proving, ; thakur2024an, ), and 300 of the 400 theorems are published after GPT-4’s training cutoff.
@@
Comparisons to State-of-the-Art and Baselines. We compare Cobblestone to 5 prior proof-synthesis tools— (1) Proverbot9001 (Sanchez20, ), an RNN-based neural theorem prover; (2) CoqHammer (Czajka18, ), which uses SMT solvers and proof reconstruction procedures; (3) Tactician (Blaauwbroek20, ), which uses an online k-nearest-neighbor algorithm to select tactics during proof search; (4) PALM (lu2024proofautomationlargelanguage, ), which leverages retrieval augmentation (RAG), custom repair prompts, and a backtracking algorithm; and (5) Rango (thompson2025rango, ), which fine-tunes an LLM model and uses RAG that includes relevant lemmas and proof scripts. Proverbot9001 outperforms other ML-based proof synthesis tools that do not use LLMs, including ASTactic (Yang19, ), TacTok (First20oopsla, ), Diva (First22icse, ), and Passport (Sanchez-Stern23toplas, ). Proverbot9001 and CoqHammer prove 19.8% and 26.6% of theorems, respectively, on CoqGym, and prove 17% and 30% on CoqGym100. In a prior “informal comparison” of tools on a subset of CoqGym, Tactician outperformed Proverbot9001 (except on one project) and sometimes outperformed CoqHammer across projects (Blaauwbroek24ICML, ).
@@
For CoqHammer, we use the Z3 (de2008z3, ), CVC4 (DBLP:conf/cav/BarrettCDHJKRT11, ), Vampire (10.1007/978-3-642-19835-9_7, ), and E (SCV:CADE-2019, ) SMT solvers, along with the default timeout settings (a 20 second prover timeout and a 5 second reconstruction timeout), following prior work (Sanchez20, ; Blaauwbroek24ICML, ). We evaluate Tactician’s ability to produce a proof fully automatically within a 10-minute timeout. Unlike Tactician, Proverbot9001 uses depth limits rather than a time limit, but usually finished its search within 10 minutes as well. We run PALM with gpt-4-0613 as its base model. One run of PALM on CoqGym100 used  3.7K tokens per theorem, 15 times less than the amount for Cobblestone (59K tokens per theorem). To correct for this, we run PALM 15 times for each dataset. Rango is run with the same setup noted in the paper—using a fine-tuned model with temperature 1.0 and a 10 minute timeout. Rango is run on an NVIDIA RTX 2080 GPU, along with a CPU with 16GB of RAM for proof checking.
@@
We also compare against 2 new baselines, ChainOfThought, and LinearRegen. ChainOfThought prompts gpt-4-0613 20 times (the same as Cobblestone) in exactly the same way as Cobblestone does (using preceding lemmas and chain-of-thought prompting as described in Section 2.4) but only checks the proofs for correctness and does no further processing. LinearRegen also prompts gpt-4-0613 up to 20 times in exactly the same way as Cobblestone, but regenerates portions of a proof starting from its first failure point. When a proof has a failure, LinearRegen keeps the prefix of the proof that succeeds, and prompts the LLM to continue from the last successful proof state.
@@
Metrics. In line with prior evaluations (First20oopsla, ; First22icse, ; First23fse, ; Sanchez-Stern23toplas, ), we use two metrics: success rate and added value. The success rate of a tool is the fraction of all theorems the tool is able to prove. The added value of tool X over tool Y is the number of new theorems X proves that Y does not, divided by the number of theorems Y proves. To measure the number of theorems a tool can prove in a fixed number of LLM samples, we define the proven@k metric. This metric is the sum of the pass@k (chen2021evaluating, ) values for each proof in a dataset, and represents the expected number of theorems proven with k samples.
@@
Computing Resources. Throughout our evaluation, we use OpenAI’s GPT-4 model (gpt-4-0613), which has a context length of 8,192 tokens. In Section 3.2, we additionally use GPT-3.5-turbo (gpt-3.5-turbo-0125), and Claude 3 Opus (claude-3-opus-
20240229
@@
). GPT-3.5-turbo has a context length of 16,385. Claude 3 Opus has a context length of 200,000 tokens, but to manage the cost of inference, we capped its token limit at 10,000. We chose these three models because their APIs support function calling, a key part of how our implementation samples whole proofs. We interface with GPT-3.5-turbo and GPT-4 through the OpenAI Python library. We use the official Anthropic library to interface with Claude 3 Opus. For each call to an LLM, we sample with temperature 1.0. All experiments are run on machines with Intel Xeon Gold 6230 CPUs and 125GB of memory. We have access to 10 cores from these CPUs, as our experiments were run in a virtualized environment.
@@ 3.2. RQ1: How Does Cobblestone Compare to State-of-the-Art Proof Synthesis Methods?
Figure LABEL:tab:rq1 shows the success rates for non-LLM tools (CoqHammer, Proverbot9001, and Tactician), our newly created baselines (ChainOfThought and LinearRegen), LLM-based tools (PALM and Rango), and Cobblestone. Note that Tactician and Rango cannot run on certain datasets (reported with N/A), as these datasets use unsupported older Coq versions.
@@
Cobblestone proves more theorems than both ChainOfThought and LinearRegen in all datasets, adding a value of up to 100.0% over ChainOfThought, and up to 104.8% over LinearRegen. Cobblestone is also more token-efficient than LinearRegen. On CoqGym100 and Wigderson100, Cobblestone uses significantly fewer tokens (59.3K vs. 99.3K on CoqGym100 and 42.0K vs. 72.3K on Wigderson100). Similarly, on PnVRocqLib100, Cobblestone uses 51.8K, while LinearRegen uses 83.7K, and on BB5100, Cobblestone uses 47.8K while LinearRegen uses 80.6K.
Cobblestone also consistently proves theorems that the other LLM-based tools cannot in all datasets, outrightly proving more theorems than these tools in many cases. Cobblestone proves more theorems than a single run of PALM (PALM 1x in Figure LABEL:tab:rq1). It also proves more theorems than our PALM benchmark (which is run 15x) on Wigderson100, PnVRocqLib100, and BB5100, proving 8%, 9% and 2% more theorems (the difference in success rates) respectively. Furthermore, in all datasets, Cobblestone proves many theorems that PALM cannot. Cobblestone proves 46 theorems PALM cannot, PALM proves 32 theorems Cobblestone cannot, and both are able to prove an overlap of 126 theorems. On CoqGym100 and Wigderson100, Cobblestone uses a similar number of tokens to PALM, with an average of 59.3K vs. PALM’s 58.1K on CoqGym100 and 42.0K vs. PALM’s 41.8K on Wigderson100. On PnVRocqLib100, Cobblestone uses fewer tokens on average (51.8K vs. PALM’s 62.0K), while using more tokens on BB5100 (47.8K vs. PALM’s 40.7K). Cobblestone and PALM’s complementarity indicate that Cobblestone may benefit from using techniques such as PALM’s retrieval augmentation (RAG) and LLM-based proof repair.
Cobblestone proves the same number of theorems as Rango on BB5100, and significantly more on PnVRocqLib100. In both cases, Cobblestone and Rango also demonstrate significant complementarity. On BB5100, their union proves 55%, and on PnVRocqLib100, it proves 47%, more than either tool alone. Across all datasets, Cobblestone proves 29 theorems that Rango cannot, Rango proves 15 theorems that Cobblestone cannot, and they both prove an overlap of 55 theorems. This indicates that Cobblestone may also benefit from leveraging Rango’s RAG. Cobblestone and Rango’s token consumption is not straightforward to compare, as Rango uses a 1.3 billion parameter LLM, which is not directly comparable to gpt-4-0613, a much larger model.
Since Cobblestone uses an LLM as a black box, it works with any LLM that is fine-tuned for tool use. We modified Cobblestone and ChainOfThought to use GPT-3.5-turbo and Claude 3 Opus, as shown in Figure 7(b), and evaluated them on CoqGym100 and Wigderson100. For all three models, Cobblestone outperforms the combination of ChainOfThought and CoqHammer. However, using different LLMs affects performance relative to GPT-4. GPT-3.5-turbo underperforms, both in ChainOfThought and in Cobblestone. Claude 3 Opus outperforms GPT-4 when used in ChainOfThought, and has mixed results in Cobblestone, slightly overperforming on CoqGym100, and underperforming on Wigderson100.
@@
(b) Figure 7b: Performance of Cobblestone and ChainOfThought implemented with different LLMs. “added value” is Cobblestone’s value added relative to each row.
RA1: Cobblestone outperforms non-LLM based proof generation methods and additional baselines. Cobblestone also performs comparably with LLM-based methods, complementing their results by proving theorems they cannot, and outperforming them in many cases. Cobblestone can be easily adapted to use LLMs other than GPT-4.
@@ 3.3. RQ2: How Much Does CoqHammer Contribute to Cobblestone’s Performance?
Figure 8 shows that CoqHammer plays an important role. On CoqGym100, Cobblestone-NoHammer (an ablated version Cobblestone) only proves 25% of theorems, while Cobblestone proves 48%. On Wigderson100, it’s 16% vs. 38%. On the two benchmarks, CoqHammer helps Cobblestone prove 96.0% and 137.5% more theorems.
Even running CoqHammer once improves ChainOfThought’s success rate—from 22% to 36% for CoqGym100, and from 17% to 31% for Wigderson100. However, Cobblestone outperforms this simpler combination of the two methods. Recall that Cobblestone uses CoqHammer in each invocation. Another simple combination is to invoke CoqHammer once for each theorem, succeeding if it or Cobblestone-NoHammer succeeds. Cobblestone also outperforms this approach (“Cobblestone-NoHammer $\cup$ CoqHammer”). On CoqGym100, Cobblestone adds 96.0% value over Cobblestone-NoHammer, compared to the simpler approach’s 52%. On Wigderson100, Cobblestone adds 137.5%, vs. the simpler approach’s 87.5%.
RA2: CoqHammer and Cobblestone are significantly complementary. CoqHammer contributes to Cobblestone’s performance, but Cobblestone outperforms CoqHammer and other simple combinations of the two. Even when run without CoqHammer, Cobblestone proves theorems other methods cannot.

### v4

@@ 2.6. Recursively Invoking Cobblestone
Through this greedy depth-first recursion, Cobblestone uses divide-and-conquer, iteratively simplifying the goal and recursing on simpler subgoals, increasing the probability of success. This way, Cobblestone breaks down the problem of proving the original theorem into smaller chunks, and leverage progress made in the previous iterations to assist the subsequent ones.
@@ 3. Evaluation
We evaluate Cobblestone on 4 datasets of theorems from open-source Coq projects. We compare Cobblestone to 5 prior tools and two LLM baselines of our own creation. Our evaluation answers 5 research questions:
@@ 3.1. Experimental Setup
Benchmarks. We construct our evaluation benchmarks from CoqGym’s test set of theorems (Yang and Deng, 2019), and from three other open-source projects—coq-wigderson (Phipathananunth, 2023), coq-bb5 (mxdys, 2025), and PnVRocqLib (PnVDiscord, [n. d.]).
@@
CoqGym is a widely used benchmark for evaluating proof-synthesis tools (Yang and Deng, 2019; First et al., 2020; First and Brun, 2022; Sanchez-Stern et al., 2023), comprised of 68,501 theorems and their associated human-written proofs across 124 projects. CoqGym’s test set consists of 26 projects with 12,161 theorems, including verdi-raft (distributed software correctness), goedel (Gödel’s 1st incompleteness theorem), and tree-automata (verified algorithms). Evaluating using this dataset allows for a more fair comparison to prior tools, and assessing tool efficacy across a diverse set of projects. However, CoqGym was released in 2019, and its projects existed on GitHub before then. Because the pre-training cutoff for GPT-4 is September 2021, GPT-4 has potentially seen CoqGym in its pre-training.
@@
To address this concern, we also evaluate using coq-wigderson (Phipathananunth, 2023), coq-bb5 (mxdys, 2025), and PnVRocqLib (PnVDiscord, [n. d.]). The first commit from each of these projects (March 2022, April 2024, and August 2024, respectively) is after the GPT-4 pre-training cutoff. coq-wigderson is a formal verification of Wigderson’s graph coloring algorithm. It consists of 174 theorems. PnVRocqLib contains 810 theorems about mathematics, including first order logic, Hilbert logic, and boolean algebra. coq-bb5 contains 1446 theorems about busy-beaver values.
@@
Because of the cost of performing inference on LLMs, full-scale evaluations on tens of thousands of theorems are impractical. In fact, the cost of just the LLM usage for our evaluation, including ablation studies, exceeded US\$8K. Instead, we create four 100-proof subsets of the four benchmark projects, sampling theorems at random. We call these subsets CoqGym100, Wigderson100, PnVRocqLib100, and BB5100. To keep the cost of ablations low, we evaluate on PnVRocqLib100 and BB5100 only for RQ1 (Section 3.2). The total size of this dataset (400 theorems), is in line with other prior work which uses LLMs for proof synthesis (Wang et al., 2024b, a; Thakur et al., 2024a), and 300 of the 400 theorems are published after GPT-4’s training cutoff.
@@
Comparisons to State-of-the-Art and Baselines. We compare Cobblestone to 5 prior proof-synthesis tools—(1) Proverbot9001 (Sanchez-Stern et al., 2020), an RNN-based neural theorem prover; (2) CoqHammer (Czajka and Kaliszyk, 2018), which uses SMT solvers and proof reconstruction procedures; (3) Tactician (Lasse et al., 2020), which uses an online k-nearest-neighbor algorithm to select tactics during proof search; (4) PALM (Lu et al., 2024), which leverages retrieval augmentation (RAG), custom repair prompts, and a backtracking algorithm; and (5) Rango (Thompson et al., 2025), which fine-tunes an LLM and uses RAG to include relevant lemmas and proof scripts. Proverbot9001 outperforms other ML-based proof synthesis tools that do not use LLMs, including ASTactic (Yang and Deng, 2019), TacTok (First et al., 2020), Diva (First and Brun, 2022), and Passport (Sanchez-Stern et al., 2023). Proverbot9001 and CoqHammer prove 19.8% and 26.6% of theorems, respectively, on CoqGym, and prove 17% and 30% on CoqGym100. In a prior “informal comparison” on a subset of CoqGym, Tactician outperformed Proverbot9001 (except on one project) and sometimes outperformed CoqHammer (Blaauwbroek et al., 2024).
@@
For CoqHammer, we use the Z3 (De Moura and Bjørner, 2008), CVC4 (Barrett et al., 2011), Vampire (Hoder et al., 2011), and E (Schulz et al., 2019) SMT solvers, along with the default timeout settings (a 20 second prover timeout and a 5 second reconstruction timeout), following prior work (Sanchez-Stern et al., 2020; Blaauwbroek et al., 2024). We evaluate Tactician’s ability to produce a proof fully automatically within a 10-minute timeout. Unlike Tactician, Proverbot9001 uses depth limits rather than a time limit, but usually finished its search within 10 minutes as well. We run PALM with gpt-4-0613 as its base model. One run of PALM on CoqGym100 used  3.7K tokens per theorem, 15 times less than the amount for Cobblestone (59K tokens per theorem). To correct for this, we run PALM 15 times for each dataset. Rango is run with the same setup noted in the paper—using a fine-tuned model with temperature 1.0 and a 10 minute timeout. Rango is run on an NVIDIA RTX 2080 GPU, along with a CPU with 16GB of RAM for proof checking. We do not compare against COPRA (Thakur et al., 2024a), another LLM-based tool, as it is too cost intensive to run in addition to our other experiments. In the COPRA paper, the authors only ran on a subset of theorems in CompCert that Proverbot9001 is not able to prove, citing budgetary constraints and high cost of evaluation.
@@
We also compare against 2 new baselines, ChainOfThought, and LinearRegen. ChainOfThought prompts gpt-4-0613 20 times (the same as Cobblestone) in exactly the same way as Cobblestone does (using preceding lemmas and chain-of-thought prompting as described in Section 2.4) but only checks the proofs for correctness without further processing. LinearRegen also prompts gpt-4-0613 up to 20 times in exactly the same way as Cobblestone, but regenerates the proof after its first failure point. When a proof has a failure, LinearRegen keeps the prefix of the proof that succeeds, and continues from the last successful proof state.
@@
Metrics. In line with prior evaluations (First et al., 2020; First and Brun, 2022; First et al., 2023; Sanchez-Stern et al., 2023), we use two metrics: success rate and added value. The success rate of a tool is the fraction of all theorems the tool is able to prove. The added value of tool X over tool Y is the number of new theorems X proves that Y does not, divided by the number of theorems Y proves. To measure the number of theorems a tool can prove in a fixed number of LLM samples, we define the proven@k metric. This metric is the sum of the pass@k (Chen et al., 2021) values for each proof in a dataset, and represents the expected number of theorems proven with k samples.
@@
Computing Resources. Throughout our evaluation, we use OpenAI’s GPT-4 model (gpt-4-0613), which has a context length of 8,192 tokens. In Section 3.2, we also use GPT-3.5-turbo (gpt-3.5-
turbo-0125
@@
) and Claude 3 Opus (claude-3-opus-20240229). GPT-3.5-turbo has a context length of 16,385. Claude 3 Opus has a context length of 200,000 tokens, but to manage inference cost, we capped its token limit at 10,000. We chose these 3 models because their APIs support function calling, a key part of our implementation. We call GPT-3.5-turbo and GPT-4 through the OpenAI Python library. We use the official Anthropic library to call Claude 3 Opus. For each call to an LLM, we sample with temperature 1.0. All experiments are run on machines with Intel Xeon Gold 6230 CPUs and 125GB of memory. We have access to 10 cores from these CPUs, as our experiments were run in a virtualized environment.
@@ 3.2. RQ1: How Does Cobblestone Compare to State-of-the-Art Proof Synthesis Methods?
Figure 7 shows success rates for non-LLM tools (CoqHammer, Proverbot9001, and Tactician), our newly created baselines (ChainOfThought and LinearRegen), LLM-based tools (PALM and Rango), and Cobblestone. Tactician and Rango cannot run on some datasets (reported with N/A) which use unsupported older Coq versions.
@@
success added success added success added success added
rate value rate value rate value rate value
CoqGym100 Wigderson100 PnVRocqLib100 BB5100
CoqHammer 30% 066.7% 27% 044.4% 23% 091.3% 27% 059.3%
Proverbot9001 17% 194.1% 10% 280.0% 12% 275.0% 23% 087.0%
Tactician N/A N/A 13% 200.0% 23% 108.7% 22% 113.6%
All non-LLM 33% 054.5% 32% 025.0% 35% 040.0% 41% 014.6%
ChainOfThought 25% 092.0% 19% 100.0% 25% 084.0% 31% 048.4%
LinearRegen 31% 061.3% 21% 104.8% 33% 048.5% 22% 104.5%
PALM 1x 34% 052.9% 16% 143.8% 16% 175.0% 23% 104.3%
PALM 52% 017.3% 30% 040.0% 35% 042.9% 41% 026.8%
PALM + 61% 42% 50% 52%
Cobblestone
Rango N/A N/A N/A N/A 30% 056.7% 43% 025.6%
Rango + N/A N/A 47% 54%
Cobblestone
Cobblestone 48% 38% 44% 43%
All Together 63% 48% 59% 61%
Figure 7. Cobblestone’s performance vs. baselines and other proof-synthesis tools. The “added value” columns report the percent of theorems Cobblestone proves over each of the other tools.
Cobblestone proves more theorems than ChainOfThought and LinearRegen in all datasets, adding value of up to 100.0% over ChainOfThought and up to 104.8% over LinearRegen. Cobblestone is also more token-efficient than LinearRegen. On CoqGym100 and Wigderson100, Cobblestone uses far fewer tokens (59.3K vs. 99.3K on CoqGym100 and 42.0K vs. 72.3K on Wigderson100). On PnVRocqLib100, Cobblestone uses 51.8K, vs. LinearRegen’s 83.7K, and on BB5100, Cobblestone uses 47.8K vs. LinearRegen’s 80.6K.
@@
Figure 8. Performance of Cobblestone and ChainOfThought implemented with different LLMs. “added value” is Cobblestone’s value added relative to each row.
@@
Cobblestone also consistently proves theorems that other LLM-based tools cannot in all datasets, outrightly proving more theorems than these tools in many cases. Cobblestone proves more theorems than a single run of PALM (PALM 1x in Figure 7). It also outperforms our PALM benchmark (run 15x) on Wigderson100, PnVRocqLib100, and BB5100, proving 8%, 9% and 2% more theorems (the difference in success rates) respectively. Furthermore, in all datasets, Cobblestone proves many theorems that PALM cannot. Cobblestone alone prove 47 theorems, PALM alone proves 32 theorems, and both can prove an overlap of 126 theorems. On CoqGym100 and Wigderson100, Cobblestone uses a similar number of tokens to PALM, with an average of 59.3K vs. PALM’s 58.1K on CoqGym100 and 42.0K vs. PALM’s 41.8K on Wigderson100. On PnVRocqLib100, Cobblestone uses fewer tokens on average (51.8K vs. PALM’s 62.0K), while using more on BB5100 (47.8K vs. PALM’s 40.7K). Cobblestone and PALM’s complementarity indicates that Cobblestone may benefit from techniques such as PALM’s retrieval augmentation (RAG) and LLM-based proof repair.
Cobblestone proves the same number of theorems as Rango on BB5100, and significantly more on PnVRocqLib100. In both cases, Cobblestone and Rango also demonstrate significant complementarity. On BB5100, their union proves 54%, and on PnVRocqLib100, it proves 47%, more than either tool alone. Across all datasets, Cobblestone alone proves 28 theorems, Rango alone proves 15, and they both prove an overlap of 55 theorems. This indicates that Cobblestone may also benefit from leveraging Rango’s RAG. Cobblestone and Rango’s token consumption is not straightforward to compare, as Rango uses a 1.3 billion parameter LLM, which is not directly comparable to gpt-4-0613, a much larger model.
Since Cobblestone uses an LLM as a black box, it works with any LLM that supports tool use. We modified Cobblestone and ChainOfThought to use GPT-3.5-turbo and Claude 3 Opus, as shown in Figure 8. For all 3 models, Cobblestone outperforms ChainOfThought $\cup$ CoqHammer. However, LLM choice can affect performance relative to GPT-4. GPT-3.5-turbo underperforms, both in ChainOfThought and in Cobblestone. Claude 3 Opus overperforms in ChainOfThought, with mixed results in Cobblestone.
[…]
Figure 9 shows that CoqHammer plays an important role. On CoqGym100, Cobblestone-NoHammer (an ablated version of Cobblestone) proves 25% of theorems, while Cobblestone proves 48%. On Wigderson100, 16% vs. 38%. On both benchmarks, CoqHammer helps Cobblestone prove 96.0% and 137.5% more theorems.
Even running CoqHammer once improves ChainOfThought’s success rate—from 22% to 36% for CoqGym100, and from 17% to 31% for Wigderson100. However, Cobblestone outperforms this simpler combination of the two. Recall that Cobblestone uses CoqHammer in each invocation. Another simple combination is to invoke CoqHammer once per theorem, succeeding if it or Cobblestone-NoHammer succeeds. Cobblestone also outperforms this approach (“Cobblestone-NoHammer $\cup$ CoqHammer”). On CoqGym100, Cobblestone adds 96.0% value over Cobblestone-NoHammer, compared to the simpler approach’s 52%. On Wigderson100, Cobblestone adds 137.5%, vs. the simpler approach’s 87.5%.
[…]